# Adversarial Robustness for Deep Learning-based Wildfire Prediction Models

### v2

@@ Abstract
Smoke detection using Deep Neural Networks (DNNs) is an effective approach for early wildfire detection. However, because smoke is temporally and spatially anomalous, there are limitations in collecting sufficient training data. This raises overfitting and bias concerns in existing DNN-based wildfire detection models. Thus, we introduce WARP (Wildfire Adversarial Robustness Procedure), the first model-agnostic framework for evaluating the adversarial robustness of DNN-based wildfire detection models. WARP addresses limitations in smoke image diversity using global and local adversarial attack methods. The global attack method uses image-contextualized Gaussian noise, while the local attack method uses patch noise injection, tailored to address critical aspects of wildfire detection. Leveraging WARP’s model-agnostic capabilities, we assess the adversarial robustness of real-time Convolutional Neural Networks (CNNs) and Transformers. The analysis revealed valuable insights into the models’ limitations. Specifically, the global attack method demonstrates that the Transformer model has more than 70% precision degradation than the CNN against global noise. In contrast, the local attack method shows that both models are susceptible to cloud image injections when detecting smoke-positive instances, suggesting a need for model improvements through data augmentation. WARP’s comprehensive robustness analysis contributed to the development of wildfire-specific data augmentation strategies, marking a step toward practicality.
Keywords wildfire  $\cdot$ smoke  $\cdot$ adversarial robustness  $\cdot$ computer vision  $\cdot$ deep learning  $\cdot$ deep neural networks  $\cdot$ noise
@@ 1 Introduction
Wildfires were among the deadliest natural disasters in 2023, surpassing storms and floods [1]. The five-year average cost of firefighting in the US is estimated at \$2.3 billion. With the number of wildfires on the rise, developing solutions for early wildfire detection is a societal issue of utmost urgency.
Several works have proposed automated solutions with a particular focus on detecting early indications of wildfires. With rapid advancements in computer vision, image-based solutions have attracted substantial attention recently. Yazdi et al. [2] conducted a comparative study of existing wildfire monitoring approaches, including high [3] and low-altitude remote sensing [4], local sensing [5], and terrestrial surveillance [2]. They concluded that using deep learning with terrestrial surveillance to detect wildfire smoke was the most effective method of predicting wildfires. Terrestrial surveillance cameras positioned at vantage points have continuous temporal coverage with a wide field of view, producing high-resolution images. This allows wildfires to be detected in their incipient stage, the earliest stage of wildfires characterized by small smoke plumes. This makes terrestrial surveillance the most competitive monitoring approach.
Wildfire smoke detection using deep learning typically involves object detection, a highly established task in computer vision, which consists of identifying and locating an object with a bounding box within an image. The target object is wildfire smoke in this context. Figure 1 illustrates wildfire smoke object detection from surveillance video images using deep learning. To develop object detection models, large-scale deep neural network (DNN) models are trained on image data usually in quantities well over the thousands for each target object. Using the trained DNN models, smoke objects can be detected in real-time, creating a powerful synergy with continuous surveillance.
@@
Figure 1: Smoke object detection from surveillance video images. A DNN object detection model creates bounding boxes to locate smoke as the target object. Image adapted from [6].
@@ 1.1 Existing DNN Solutions to Wildfire Smoke Detection
Existing wildfire smoke detection frameworks in the literature are characterized by two major DNN architectures: CNN and Transformer. For the CNN-based approaches, Jeong et al. [7] proposed a real-time wildfire detection model, where a YOLOv3 [8]-based CNN framework (for producing candidate proposals for smoke-positive regions) is combined with Long-Short Term Memory (LSTM) [9] (for screening the candidates). While enforcing temporal consistency is demonstrated to be effective in distinguishing smoke and cloud, the model was trained on samples taken from only 12 scenes. Despite a large number of image frames, the effective sample size may be as few as 12, potentially limiting the model’s ability to generalize. Additionally, for wildfire flame detection, which is considerably easier than smoke detection, Al-Smadi et al. [10] developed another CNN-based detection model using newer version of YOLOs to report on extremely high prediction performance. However, the model was trained on positive-only samples with a relatively small image count (1,723). It is not entirely clear how the limited sample size and synthetically inflated wildfire occurrences affect model robustness and generalizability.
For the transformer-based approach, there has only been one work in the literature so far. Yazdi et al. [2] proposed NEMO (NEvada sMOke detection benchmark), using the DEtection TRansformer (DETR) [11] framework for the first time in wildfire smoke detection. They set the practical benchmark for state-of-the-art precision (See Table 3). To address the issue of false positives, they use several data preprocessing strategies to artificially diversify data. Specifically, they inject 260 smoke-negative background images randomly selected from the internet. They also create 116 collages from reused smoke-positive and smoke-negative images to further address this problem. Generally, collages were effective not only for reducing false positives but for introducing variety in the object’s position and size. It was shown that this preprocessing step can reduce the false alarm rate. However, detailed discussions on the model’s generalizability and robustness on real-life data are not present beyond an ablation study of the detection encoder mechanism of the model and a typical time-series detection analysis.
@@ 1.2 Challenges in Adapting DNN Solutions to Wildfire Smoke Detection
While the DNN-based object detection approach has shown promising results [10, 7, 2], two main challenges remain: 1) the limited sample size coupled with low image diversity in wildfire smoke training images, and 2) the lack of universal solutions for improving model robustness.
The first challenge arises because of the properties of wildfire smoke. Smoke is spatially anomalous because it occupies a small portion relative to the entire image. As a result, it is absent from the vast majority of pixels in the frames of high-resolution continuous surveillance videos. This makes manual annotation of images for object detection highly labor-intensive, necessitating ad-hoc preprocessing steps as discussed in Section 1.1. Furthermore, smoke is temporally anomalous becuaise it originates from it being a rare event. This results in surveillance image sequences having relatively small smoke-positive frames compared to the entire data. When trained on limited datasets, DNNs are known to produce unexpected and/or detrimental outcomes even from slight modifications to the input [12]. In wildfire smoke detection, where training data are almost always limited, this problem can be exacerbated by the current trend of shifting from CNNs (Convolutional Neural Networks) to Transformer-based models [13]. Despite the importance of this issue, little work has been done to assess the robustness of wildfire detection models across various inputs.
The second challenge stems from the black-box nature of DNNs. Since the internal workings of the model, including feature extraction processes and the role of parameters, are not immediately discernible, addressing potential issues requires significant case-by-case effort, even if the source code is available. The architecture of CNNs and Transformers differs greatly, making analysis based on model-specific quantities (e.g., gradients) less practical. Since the spatiotemporal anomaly issue could potentially introduce vulnerabilities to the model, a model-agnostic framework for evaluating the model’s robustness and fine-tuning must be developed.
@@ 1.3 Contributions
To address these challenges, we propose WARP (Wildfire Adversarial Robustness Procedure), the first model-agnostic framework to comprehensively evaluate the adversarial robustness of wildfire detection models. Unlike common adversarial attack methods, which often rely on solving an optimization problem for perturbations using internal model details, WARP is model-agnostic and uses relatively simple noise injection methods. Additionally, WARP incorporates wildfire-specific contexts in designing noise tests, especially in the distinction between smoke and cloud, rather than using generic random noise. To the best of our knowledge, WARP is the first framework to offer a model-agnostic, contextual adversarial robustness evaluation method specifically tailored for wildfire detection. Insights obtained through WARP’s analysis can be used to further improve the model through data augmentation, making it a crucial first step toward a truly practical/reliable wildfire detection system. The main contributions of our study are summarized as follows.
@@ 2.1 Problem Statement
We assume that an image-based wildfire detection model $y=f(x)$ is given, where $x$ is an input image, typically an image frame captured by a surveillance camera, and $y\in\{0,1\}$ is the binary label indicating whether $x$ contains wildfire smoke ($y=1$) or not ($y=0$). Since wildfire smoke is spatially localized, the function $f$ incorporates two steps: bounding box generation (locating subimages within $x$) and scoring (estimating the probability of containing smoke), as illustrated in Fig. 1.
Additionally, we assume that a test dataset $\mathcal{D}=\{x^{(1)},\ldots,x^{(N)}\}$ is available, where $x^{(n)}$ is the $n$-th image. Since the input images are often noisy due to varying weather and lighting conditions, a practical wildfire detection system must be robust to small variations in test images. The central question we address in this papers is how robust such a wildfire detection model is.
We address this question by proposing metrics indicating changes of the classification outcomes upon introducing small perturbations to the input image. Specifically, as illustrated in Fig. 2, we compute the precision degradation metric $L$, the classification flip probabilities $\alpha_{i}$ and $\beta_{i}$, and the localization deception rate $\gamma_{i}$. $L$ (see Section 3.1) seeks to quantify the model’s robustness against global noise, whereas $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ (see Section 3.2 seek to quantify the model’s robustness against local noise.
As discussed, wildfire smoke detection models need an adaptable and contextualized adversarial robustness evaluation method. For a disaster prevention task where human lives are involved, wildfire detection solutions must be robust. That is, a DNN wildfire prediction model must be able to adapt to a variety of wildfire scenarios, not exclusive to the wildfires it has seen in its training data. We leverage adversarial attacks to identify any critical model vulnerabilities through a variety of noise perturbations. Insights obtained will be vital especially in a task like image-based wildfire smoke detection, as they can be used to further diversify the training data through data augmentation which will be discussed in Section 5.
@@ 2.2 Adversarial Robustness
Adversarial robustness refers to a model’s resilience to external attacks on its input. “Attacks" usually involve the injection of noise perturbations (slight modifications in the form of noise) onto the original input to artificially create counterfactual (unforeseen) test scenarios. These are called adversarial examples.
@@
If $f(x)$ matched the ground truth, but $f(x+r)$ does not, the model cannot be considered robust. However, it is conditionally required that $r$ be reasonably small, as it is almost guaranteed that the classification output will change if $r$ is arbitrarily large. Therefore, the most desirable perturbed image tensor $x^{\prime}\triangleq x+r$ looks almost identical to the original image tensor $x$, but has the maximum impact on classification outcome. The optimization problem to find such a perturbation is called an adversarial attack [14]:
@@
Depending on their nature, perturbations can generally be categorized into two categories: global and local attacks [15]. 1) Global Attacks: Both [16] and [12] proposed global attack methods that seek the optimal perturbation that covers the entire input image. [16] incorporated the distant constraint as a penalty term $c\|r\|_{1}$, where $c$ is a constant and $\|\cdot\|_{1}$ is the $\ell_{1}$ norm, that is added to the objective function. Moreover, [12] proposed a more computationally efficient approach by constraining $r$ not by a constant but only to the sign of the gradient of the loss function $\frac{\partial}{\partial x}\mathrm{Loss}(x,y)$. 2) Local Attacks: Both [17] and [18] proposed local attack methods. [18] proposed a single-pixel attack method based on a differential evolution framework. However, single-pixel attacks are highly inefficient for the typical object detection input which is a $640\times 640$ pixel image. Additionally, [17] proposed a patch-based method, where a constant patch would be directly injected into the image. These patches adapt to different backgrounds and transformations, including position, size, and rotation. Patches were generally successful at deceiving otherwise rigorously trained models, thus making them a prime tool for adversarial robustness.
@@
Moreover, local attack methods typically disregard the context of their task. Both single-pixel attacks and patch noise attacks are generally abstract formulations when viewed by humans, and generally do not fit in the context of wildfire smoke detection (i.e., abstract glob-like patches will not appear in camera surveillance images). Thus, their utility may be limited to deceiving the model in highly specific conditions. While this kind of noise fine-tuning can be useful for models trained to detect everyday objects on massive quantities of data, wildfire smoke detection is a more nuanced task. Contextual cues, such as the subtle differences in objects’ shape, color, and position, are paramount to distinguish smoke from clouds under limited data. Therefore, contextual awareness is essential for developing robust and reliable smoke detection systems.

### v3

@@ Abstract
Rapidly growing wildfires have recently devastated societal assets, exposing a critical need for early warning systems to expedite relief efforts. Smoke detection using camera-based Deep Neural Networks (DNNs) offers a promising solution for wildfire prediction. However, the rarity of smoke across time and space limits training data, raising model overfitting and bias concerns. Current DNNs, primarily Convolutional Neural Networks (CNNs) and transformers, complicate robustness evaluation due to architectural differences. To address these challenges, we introduce WARP (Wildfire Adversarial Robustness Procedure), the first model-agnostic framework for evaluating wildfire detection models’ adversarial robustness. WARP addresses inherent limitations in data diversity by generating adversarial examples through image-global and -local perturbations. Global and local attacks superimpose Gaussian noise and PNG patches onto image inputs, respectively; this suits both CNNs and transformers while generating realistic adversarial scenarios. Using WARP, we assessed real-time CNNs and Transformers, uncovering key vulnerabilities. At times, transformers exhibited over 70% precision degradation under global attacks, while both models generally struggled to differentiate cloud-like PNG patches from real smoke during local attacks. To enhance model robustness, we proposed four wildfire-oriented data augmentation techniques based on WARP’s methodology and results, which diversify smoke image data and improve model precision and robustness. These advancements represent a substantial step toward developing a reliable early wildfire warning system, which may be our first safeguard against wildfire destruction.
Keywords wildfire  $\cdot$ smoke  $\cdot$ adversarial robustness  $\cdot$ computer vision  $\cdot$ deep neural networks  $\cdot$ noise
@@ 1 Introduction
Wildfires were among the deadliest US natural disasters in 2023, surpassing storms and floods [30]. They are also some of the costliest, with the US five-year average cost of firefighting estimated at USD 2.3 billion. Indeed, recent wildfires such as the 2025 Palisades Fires have demonstrated that uncontained fires may have a substantial cost on human life and infrastructure even in a short time frame [34]. As wildfires become more frequent, early wildfire detection systems are imperative for evacuation and prevention.
Several works have proposed automated solutions with a particular focus on detecting early indications of wildfires. With rapid advancements in computer vision, image-based solutions have attracted substantial attention recently. Yazdi et al. [39] conducted a comparative study of existing wildfire monitoring approaches, including high- [13] and low-altitude remote sensing [17], local sensing [3], and terrestrial surveillance [39]. They concluded that using deep learning with terrestrial surveillance to detect wildfire smoke was the most effective method of predicting wildfires. Terrestrial surveillance cameras positioned at vantage points have continuous temporal coverage with a wide field of view, producing high-resolution images. This allows wildfires to be detected in their incipient stage, the earliest stage of wildfires which is characterized by small smoke plumes. This makes terrestrial surveillance the most competitive monitoring approach.
Wildfire smoke detection using deep learning typically involves object detection, a highly established task in computer vision. This consists of identifying and locating an object with a bounding box within an image. The target object is wildfire smoke in this context. Figure 1 illustrates wildfire smoke object detection from surveillance video images using deep learning. To develop object detection models, large-scale deep neural network (DNN) models are trained on image data, usually in quantities well over the thousands for each target object. Using the trained DNN models, smoke objects can be detected in real time, creating a powerful synergy with continuous surveillance. By doing so, automatic prediction systems can be developed which can (1) replace human resources that are expensive and time-consuming, and (2) potentially detect with a higher precision than their human counterparts. In tasks like wildfire detection, which is both costly [27] and demands accuracy (see Section 1.2), these advantages of DNNs are particularly desirable.
@@
Figure 1: Smoke object detection from surveillance video images. A DNN object detection model creates bounding boxes (see green box) to locate smoke as the target object. Image adapted from [2].
@@ 1.1 Existing DNN Solutions to Wildfire Smoke Detection
Existing camera-based wildfire smoke detection using computer vision frameworks in the literature are characterized by two major DNN architectures: CNN and transformer. There are numerous CNN-based approaches [15, 23, 22, 12, 1]; in particular, Jeong et al. [22] proposed a real-time wildfire detection model, where a YOLOv3 [29]-based CNN framework (for producing candidate proposals for smoke-positive regions) is combined with Long-Short Term Memory (LSTM) [20] (for screening the candidates). While enforcing temporal consistency is demonstrated to be effective in distinguishing smoke and cloud, the model was trained on samples taken from only 12 scenes. Despite a large number of image frames, the effective sample size may be as few as 12, potentially limiting the model’s ability to generalize. Additionally, for wildfire flame detection, which is considerably easier than smoke detection, Al-Smadi et al. [1] developed another CNN-based detection model using newer versions of YOLO to report an extremely high prediction performance. However, the model was trained on positive-only samples with a relatively small image count (1723). It is not entirely clear how the limited sample size and synthetically inflated wildfire occurrences affect model robustness and generalizability.
For the transformer approach, there have only been two camera-based transformer approaches to wildfire detection in the literature, to the best of our knowledge [39, 37].  In particular, Yazdi et al. [39] proposed NEMO (NEvada sMOke detection benchmark), using the DEtection TRansformer (DETR) [7] framework for the first time in wildfire smoke detection. They set the practical benchmark for state-of-the-art precision. To address the issue of false positives, they use several data preprocessing strategies to artificially diversify data. Specifically, they inject 260 smoke-negative background images randomly selected from the internet. They also create 116 collages from reused smoke-positive and smoke-negative images to further address this problem. Generally, the collages were effective not only for reducing false positives but for introducing variety to the object’s position and size. It was shown that this preprocessing step can reduce the false alarm rate. However, detailed discussions on the model’s generalizability and robustness on real-life data are not present beyond an ablation study of the model’s detection encoder mechanism and a typical time-series detection analysis.
@@ 1.2 Challenges in Adapting DNN Solutions to Wildfire Smoke Detection
While the DNN-based object detection approach has shown promising results [1, 22, 39], two main challenges remain: (1) the limited sample size coupled with low diversity in wildfire smoke training images, and (2) the lack of universal solutions for improving model robustness.
The first challenge arises because of the properties of wildfire smoke. Smoke is spatially anomalous because it occupies a small portion relative to the entire image. As a result, it is absent from the vast majority of pixels in the frames of high-resolution continuous surveillance videos. This makes manual annotation of images for object detection highly labor-intensive, necessitating ad hoc preprocessing steps as discussed in Section 1.1. Furthermore, smoke is temporally anomalous because it originates from a rare event. This results in surveillance image sequences having relatively small smoke-positive frames compared to the entire data. Thus, limited data generally create imbalance problems, [9] and when used for training, can cause detrimental performance degradation [6, 24]. Furthermore, when trained on limited datasets, DNNs are known to produce unexpected and/or detrimental outcomes even from slight modifications to the input [16]. In wildfire smoke detection, where training data are almost always limited, this problem can be exacerbated by the current trend of shifting from CNNs (Convolutional Neural Networks) to transformer-based models [19], as the latter explicitly handles second-order statistics through the key-value transformation  [11]. Despite the importance of this issue, little work has been done to assess the robustness of wildfire detection models across various inputs.
The second challenge arises because DNNs are inherently black boxes [26]. Since the internal workings of the model, including feature extraction processes and the role of parameters, are not immediately discernible, addressing potential issues requires substantial case-by-case effort, even if the source code is available. The architecture of CNNs and transformers differs greatly, making analysis based on model-specific quantities (e.g., gradients) less practical. Since the spatiotemporal anomaly issue could potentially introduce vulnerabilities to the model, a model-agnostic framework for evaluating the model’s robustness and fine-tuning must be developed.
@@ 1.3 Contributions
We recognize the following advantages of DNNs for wildfire detection: previous studies have found that DNN–computer vision models are the most effective in wildfire detection due to their compatibility with surveillance cameras, which have the most coverage compared to other wildfire observation methods. They also synergize well with cameras because they can keep up with their footage in real time, allowing for an automated wildfire detection system that saves costs and human errors.
However, we also recognize the following disadvantages of DNNs for wildfire detection: because of smoke’s fundamental properties, smoke image data is highly difficult to collect, which not only creates a severe data shortage but also a class imbalance problem. This limits the model’s ability to be generalized to all wildfire scenarios. More importantly, DNNs are black boxes, meaning their internal decision-making mechanisms are too abstract to interpret. This makes it challenging to identify specific sources of robustness vulnerabilities. Not only that, these mechanisms differ greatly across architectures, making comparing robustness across different models difficult.
To address these challenges, we propose WARP (Wildfire Adversarial Robustness Procedure), the first model-agnostic framework to comprehensively evaluate the adversarial robustness of wildfire detection models. Unlike common adversarial attack methods, which often rely on solving an optimization problem for perturbations using internal model details, WARP is model-agnostic and uses relatively simple noise injection methods. Additionally, WARP incorporates wildfire-specific contexts in designing noise tests, especially in the distinction between smoke and cloud, rather than using generic random noise. To the best of our knowledge, WARP is the first framework to offer a model-agnostic, contextual adversarial robustness evaluation method specifically tailored for wildfire detection. Insights obtained through WARP’s analysis can be used to further improve the model through data augmentation, making it a crucial first step toward a truly practical/reliable wildfire detection system. The main contributions of our study are summarized as follows:
@@ 2.1 Problem Statement
We assume that an image-based wildfire detection model $y=f(x)$ is given, where $x$ is an input image, typically an image frame captured by a surveillance camera, and $y\in\{0,1\}$ is the binary label indicating whether $x$ contains wildfire smoke ($y=1$) or not ($y=0$). Since wildfire smoke is spatially localized, the function $f$ incorporates two steps: bounding box generation (locating subimages within $x$) and scoring (estimating the probability of containing smoke), as illustrated in Figure 1.
Additionally, we assume that a test dataset $\mathcal{D}=\{x^{(1)},\ldots,x^{(N)}\}$ is available, where $x^{(n)}$ is the $n$-th image. Since the input images are often noisy due to varying weather and lighting conditions, a practical wildfire detection system must be robust to small variations in test images. The central question we address in this study is how robust such a wildfire detection model is.
We address this question by proposing metrics indicating changes in the classification outcomes upon introducing small perturbations to the input image. Specifically, as illustrated in Figure 2, we compute the precision degradation metric $L$, the classification flip probabilities $\alpha_{i}$ and $\beta_{i}$, and the localization deception rate $\gamma_{i}$. $L$ (see Section 3.1) seeks to quantify the model’s robustness against global noise, whereas $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ (see Section 3.2 seek to quantify the model’s robustness against local noise.
Figure 2: WARP workflow.
As discussed, wildfire smoke detection models need an adaptable and contextualized adversarial robustness evaluation method. For a disaster prevention task where human lives are involved, wildfire detection solutions must be robust. That is, a DNN wildfire prediction model must be able to adapt to a variety of wildfire scenarios, not exclusive to the wildfires it has seen in its training data. We leverage adversarial attacks to identify any critical model vulnerabilities through a variety of noise perturbations. The insights obtained will be vital, especially in a task like image-based wildfire smoke detection, as they can be used to further diversify the training data through data augmentation which will be discussed in Section 5.
@@ 2.2 Adversarial Robustness
Adversarial robustness refers to a model’s resilience to external attacks on its input. “Attacks” usually involve the injection of noise perturbations (slight modifications in the form of noise) into the original input to artificially create counterfactual (unforeseen) test scenarios. These are called adversarial examples.
@@
If $f(x)$ matches the ground truth, but $f(x+r)$ does not, the model cannot be considered robust. However, it is conditionally required that $r$ be reasonably small, as it is almost guaranteed that the classification output will change if $r$ is arbitrarily large. Therefore, the most desirable perturbed image tensor $x^{\prime}\triangleq x+r$ looks almost identical to the original image tensor $x$, but has the maximum impact on classification outcome. The optimization problem to find such a perturbation is called an adversarial attack [26]:
@@
Depending on their nature, perturbations can generally be categorized into two categories: global and local attacks [8]. (1) Global Attacks: The authors of [33, 16] proposed global attack methods that seek the optimal perturbation that covers the entire input image. The authors of [33] incorporated the distant constraint as a penalty term $c\|r\|_{1}$, where $c$ is a constant and $\|\cdot\|_{1}$ is the $\ell_{1}$ norm, that is added to the objective function. Moreover, [16] proposed a more computationally efficient approach by constraining $r$ not by a constant but only to the sign of the gradient of the loss function $\frac{\partial}{\partial x}\mathrm{Loss}(x,y)$. (2) Local Attacks: The authors of [5, 32] proposed local attack methods. [32] proposed a single-pixel attack method based on a differential evolution framework. However, single-pixel attacks are highly inefficient for the typical object detection input which is a $640\times 640$ pixel image. Additionally, [5] proposed a patch-based method, where a constant patch would be directly injected into the image. These patches adapt to different backgrounds and transformations, including position, size, and rotation. Patches were generally successful at deceiving otherwise rigorously trained models, thus making them a prime tool for adversarial robustness.
@@
Moreover, local attack methods typically disregard the context of their task. Both single-pixel attacks and patch noise attacks are generally abstract formulations when viewed by humans, and generally do not fit in the context of wildfire smoke detection (i.e., abstract glob-like patches will not appear in camera surveillance images). Thus, their utility may be limited to deceiving hyper-tuned models in highly specific conditions. Wildfire smoke detection demands a more nuanced usage for noise patches. Contextual cues, such as the subtle differences in objects’ shape, color, and position, are paramount to distinguish smoke from highly similar objects (e.g., clouds, man-made structures with similar coloration) at a considerable distance. With limited wildfire data, it is not an option to make models learn these contextual cues by retraining them on additional images. Therefore, patch noise attack methods must take context into account to (1) evaluate a model’s ability to differentiate objects using the task’s context and (2) create adversarial examples that train this ability.